OpenStudy (freckles):

\[\text{ The function } f(n) \text{ is defined for all positive integers } \\ n \text{ and takes on non-negative integral values } \\ \text{ Also, for all } m, n \\ f(m+n)-f(m)-f(n)=0 \text{ or } 1 \\ f(2)=0, f(3)>0, \text{ and } f(9999)=3333. \\ \text{ Determine } f(1982).\]

OpenStudy (freckles):

under harder problems number 8 last page

OpenStudy (freckles):

@AlexandervonHumboldt2 can you please tell me what integral value means?

OpenStudy (freckles):

if you know

OpenStudy (alexandervonhumboldt2):

sorry @freckles , never heard of such term.

OpenStudy (freckles):

its k

OpenStudy (freckles):

sounds like it is just an integer

OpenStudy (freckles):

I guess I will assume it is an integer until I run into problems :p

OpenStudy (freckles):

hmmm... well I know since f(2)=0 we have \[f(2-n)+f(n)=0 \text{ or } -1 \]

OpenStudy (freckles):

and if I used f(2)=0 again... like replace n with 2 I get \[f(2-2)+f(2)=0 \text{ or } -1 \\ f(0)+f(2)=0 \text{ or } -1 \\ \text{ and this means } f(0)+0=0 \text{ or } -1 \text{ so we have} f(0)=0 \text{ or } -1 \]

imqwerty (imqwerty):

shuldn't it be 0 or +1?

OpenStudy (freckles):

well I had multiplied both sides by -1 earlier

OpenStudy (freckles):

I didn't show that but I could show you what I did

imqwerty (imqwerty):

oh okay

OpenStudy (freckles):

\[f(m+n)-f(m)-f(n)=0 \text{ or } 1 \\ \text{ replace } m+n=2 \text{ and so } m=2-n \\ f(2)-f(2-n)-f(n)=0 \text{ or } 1 \\ -f(2-n)-f(n)=0 \text{ or } 1 \] and then I multiplied both sides by -1

OpenStudy (freckles):

getting that one equation I had above

imqwerty (imqwerty):

f(1) will be 0

OpenStudy (freckles):

I got 0 or -1/2 for f(1)

OpenStudy (freckles):

\[f(2-n)+f(n) =0 \text{or } -1 \\ \text{ replace } n \text{ with } 1 \\ \\ f(1)+f(1)=0 \text{ or } -1 \\ f(1)=0 \text{ or } \frac{-1}{2} \\ \]

imqwerty (imqwerty):

but it says that the function is defined for all positive integers n and takes NON NEGATIVE values so f(1) can't be -1/2

OpenStudy (freckles):

oh I thought that meant we couldn't plug in negative integers

OpenStudy (freckles):

I didn't know it meant we couldn't have negative outputs

OpenStudy (freckles):

so if that is the case then f(0) is also 0

OpenStudy (freckles):

though I guess we couldn't use that really because the input ended up being 0 instead of positive integer

OpenStudy (freckles):

that was above when I replaced n with 2

OpenStudy (freckles):

so f(3)=1

OpenStudy (freckles):

oh my

OpenStudy (freckles):

I wonder if this is fibnooci sequence

imqwerty (imqwerty):

we know that the function will give positive value we can prove that n can't be negative \[f(3-1)-f(3)-f(-1)=0or1\]\[0-ve-f(-1)=0or1\]to make LHS 1 or 0 -f(-1) has to be positive so f(-1) has to be negative and this can't happen so we can't have negative n hmm is this special case or general

OpenStudy (freckles):

definitely not Fibonacci sequence I don't think the 9999th term is 3333

OpenStudy (freckles):

and I got f(3)=1 by the following: \[f(2+1)-f(2)-f(1)=0 \text{ or } 1 \\ f(3)-0-0=0 \text{ or } 1 \\ f(3)=1 \text{ since } f(3)>0\]

imqwerty (imqwerty):

yes

OpenStudy (freckles):

\[f(2+2)-f(2)-f(2)=0 \text{ or } 1 \\ f(4)=0 \text{ or } 1 \] we can't determine f(4) yet

OpenStudy (freckles):

actually I think I just determine it was 1

OpenStudy (freckles):

watch this... \[f(3+1)-f(3)-f(1)=0 \text{ or } 1 \\ f(4)-1-0=0 \text{ or } 1 \\ f(4)=1 \text{ or } 2 \]

imqwerty (imqwerty):

yes :)

OpenStudy (freckles):

\[f(4)=(1 \text{ or } 2 ) \text{ and } (1 \text{ or } 0)=1\]

OpenStudy (freckles):

that f(9999)=3333 seems interesting...

imqwerty (imqwerty):

tricky and confusing..nice question

OpenStudy (freckles):

yeah @Kainui kinda brought up functional equations and I said I was going to do my own research thought it would be nice to post this question since it did look tricky and fun

OpenStudy (freckles):

maybe we can try to use that cauchy equation thing on page 8

OpenStudy (freckles):

but we have a problem for the or 1 part

OpenStudy (freckles):

\[\text{ if } f(m+n)-f(m)-f(n)=0 \text{ then } f(m+n)=f(m)+f(n) \\ \text{ then } f(x)=cx \text{ where } c \text{ is a constant }\]

OpenStudy (freckles):

\[\text{ if } f(m+n)-f(m)-f(n)=1 \text{ then } f(m+n)=f(m)+f(n)+1 \\ \text{ and if } f(x)=cx+b \\ \text{ then we have } \\ c(m+n)+b=cm+b+cn+b+1 \\ c(m+n)+b=c(m+n)+2b+1 \\ b=2b+1 \\ -b=1 \\ b=-1 \] so I guess f(x)=cx-1 ...if I did that right

OpenStudy (freckles):

\[f(2)=0 \\ \text{ so } f(x)=cx \text{ or } f(x)=cx-1 \\ \text{ if } f(x)=cx \text{ then } f(2)=2c=0 \implies c=0 \text{ so } f(x)=0 \\ \text{ or we could have } \\ f(x)=cx-1 \text{ then } f(2)=2c-1=0 \implies c=\frac{1}{2} \\ \text{ and so } f(x)=\frac{1}{2}x-1\]

OpenStudy (freckles):

now f(3)>0 \[\text{ but we got } f(3)=1 \text{ and neither one fits }\]

OpenStudy (freckles):

but maybe for some values we have f(x)=0 while for other values we have f(x)=cx-1 but I solved f(3)=c(3)-1 for c and got c=2/3 and this does not work for f(9999)=3333

OpenStudy (freckles):

maybe f(x)=cx-1 isn't the most general form for f(m+n)=f(m)+f(n)+1 that could be another possibility

imqwerty (imqwerty):

f(5) will be 2

imqwerty (imqwerty):

shuld we find values like this or do some other approach

OpenStudy (freckles):

I was trying another approach

OpenStudy (freckles):

any approach you can think of

OpenStudy (freckles):

how did you get f(5)=2?

imqwerty (imqwerty):

\[f(4+1)-f(4)-f(1)=0or1\]\[f(5)-1=0or1\] \[f(2+3)-f(2)-f(3)=0or1\]\[f(5)-2=0or1\] If the 1st eq has value 0 then f(5) will bw 1 and if this happens the 2nd equation will become -1 so f(5)=1 is rejected now we have only 1 option left in 1st eq so f(5)-1=1 so f(5)=2

OpenStudy (freckles):

but f(2)=0 and f(3)=1 so -f(2)-f(3)=-0-1=-1 not -2

imqwerty (imqwerty):

): sry i made a mistake out there

OpenStudy (freckles):

don't be sad

OpenStudy (freckles):

I did find a answer. I don't know how much I like it.

OpenStudy (freckles):

I guess the integral part of the question was to be a big hint

OpenStudy (freckles):

it looks like I could of half used cauchy if I used the right condition... \[f(x)=cx \\ f(9999)=c \cdot 9999 \\ 3333=c \cdot 9999 \\ c=\frac{1}{3} \\ \text{ so } f(x)=\frac{1}{3} x\] and then I guess we were suppose to observe a pattern and say hey let's take the integer part of that answer

imqwerty (imqwerty):

yea

OpenStudy (freckles):

\[f(x)=\text{integer part of } (\frac{1}{3} x)\]

OpenStudy (freckles):

thank you imqwerty for putting much effort into this problem with me

imqwerty (imqwerty):

haha :D no problem

OpenStudy (kainui):

I'm still looking at this, I have several ideas to share however this morning has been busy, I went out to eat breakfast and planning a winter vacation with a friend in Europe so I haven't been able to chip in yet!

OpenStudy (kainui):

I used @freckles trick of changing m=2 and go this form, also I changed that 0 or 1 into a function r(n) just for my own purposes it's not really anything too special. So I start from here: \[f(k+2)-f(k)=r(k)\] From here I telescope over this sum keeping in mind that k either goes over only even or only odd numbers depending on where I start from: (If you have questions or doubts I can try to derive this, but it sort of makes more sense to me to just write it out as this) \[f(n) - f(a) = \sum_{k=0}^{(n-a-2)/2} r(2k+a)\] So I try plugging in: \[f(1982) = \sum_{k=0}^{989} r(2k+2)\] since r(n) is either 0 or 1, the best my method here does so far is puts a bound on f(1982) somewhere between 1 and 990. What remains to solve for is the general form for just the even terms of r(n). In a way though, that seems like most of the problem, all I've done is just rewrite it a little bit. Hmmm...